OpenStudy (anonymous):

Using the standard enthalpies of formation, what is the standard enthalpy of reaction? CO(g) + H2O ---> CO2(g) + H2(g) CO2= -393.5 CO= -110.5 H2= 0 H2O= -285.8 I keep getting 2.8 put it says the answer is wrong

OpenStudy (aaronq):

\(\Delta H=(-393.5)-((-285.8)+(-393.5))\) \(\Delta H=(-393.5)-(-679.3)\) \(\Delta H=-393.5+679.3)=285.8\)

OpenStudy (anonymous):

my hw saying its incorrect though

OpenStudy (aaronq):

lol oh i copied down the value for CO incorrectly \(\Delta H=(-393.5)-((-110.5)+(-285.8))=2.8 \) i get the same thing. were you given those values or did you look them up?

OpenStudy (anonymous):

i looked them up from my text book

OpenStudy (anonymous):

but thanks anyways

OpenStudy (aaronq):

for water, was it liquid or gas?

OpenStudy (anonymous):

i made the mistake by putting liquids instead of gas oopppps lol thanks

OpenStudy (aaronq):

hah there we go. no probs